# Seeing Far and Clearly: Mitigating Hallucinations in MLLMs with Attention Causal Decoding

### v1

@@ 4.1 Upper Triangular Metric as Attention Registers
To alleviate attention collapse issues, we propose the dedicated attention register to allocate excess attention scores. Given the original attention score matrix $\omega\in\mathbb{R}^{n\times n}$, we construct an upper triangular score matrix $\mathcal{P}\in\mathbb{R}^{n\times n}$ as attention register, defined as follows:
$\mathcal{P}_{ij}=\begin{cases}0,&\text{if }j\leq i,\\ -\sigma\,(j-i),&\text{if }j>i,\end{cases}$ (6)
where $\sigma$ is a decay hyperparameter. Consequently, for each row $i$, the first $i$ entries are zero(thus preserving the causal structure), and the remaining entries decay linearly with respect to $(j-i)$. To incorporate this register into the attention mechanism, we define a binary lower-triangular mask $C\in\mathbb{R}^{n\times n}$ as
$C_{ij}=\begin{cases}1,&\text{if }j\leq i,\\ 0,&\text{if }j>i.\end{cases}$ (7)
As illustrated in Fig. 4, the original attention score matrix $\omega$ is first modulated by the lower-triangular mask $C$ via element-wise multiplication, ensuring that only the current and preceding tokens contribute to the attention computation. The resulting scores are then combined with the register matrix $\mathcal{P}$ to yield the adjusted score matrix, expressed as $\mathbf{W}=(\omega\cdot C)+\mathcal{P}$, where $\cdot$ denotes element-wise multiplication. This formulation guarantees that the register $\mathcal{P}$ effectively absorbs any excess attention allocated to tokens beyond the current position, thereby preserving the essential contextual information for robust attention propagation.
Finally, to maintain the autoregressive property during decoding, the adjusted scores $\mathbf{W}$ are passed through a softmax function and subsequently masked by $C$:
$\tilde{\mathbf{W}}=\text{SoftMax}(\mathbf{W})\cdot C.$ (8)
This final operation guarantees that any attention weights assigned to future tokens are nullified, thereby preserving causal decoding.

### v2

@@ 4.1 Upper Triangular Metric as Attention Registers
To alleviate attention collapse issues, we propose the dedicated attention register to allocate excess attention scores. For each $\omega$, we construct an upper triangular score matrix $\mathcal{P}\in\mathbb{R}^{n\times n}$ as attention register, defined as follows:
$\mathcal{P}_{i}=[\underbrace{0,0,\cdots,0,}_{i}\underbrace{\mathcal{P}_{i,i+1},\mathcal{P}_{i,i+2},\cdots,\mathcal{P}_{i,n}}_{n-i}]_{n},\vspace{-0.2cm}$ (6)
where $\mathcal{P}_{i}$ allocates $n-i$ register-attention scores in each row to handle excess attention values while maintaining zero values for positions up to $i$. To integrate $\mathcal{P}$ with $\omega$, we adjust $\omega$ by adding the register-attention scores from $\mathcal{P}$ as follows:
$\mathbf{W}=\omega\cdot C+\mathcal{P},\quad\text{where}\quad C=\text{tril}(\mathbbm{1}_{n\times n}),$ (7)
where $C\in\mathbb{R}^{n\times n}$ denotes a lower-triangular matrix filled with ones to ensure causal masking by allowing attention only to preceding or current tokens, as illustrated in Fig. 4.
Since the model is training-free, the attention-registers $\mathcal{P}$ should not interfere with the original attention score distribution $\omega$ during inference and align with the relative positional encoding $R$ in Eq. 5 to maintain coherence in generated text. For $\mathcal{P}_{i,j}$, the values are defined as:
$\mathcal{P}_{i,j}=-(j-i)\cdot\sigma,\quad\forall j>i,$ (8)
where $\sigma$ is a decay rate hyperparameter. This setup ensures that $\mathcal{P}$ conforms to the gradual attenuation pattern in attention. Thus, the final attention score matrix with FarSight is defined as:
$\mathbf{W}_{i}=[\mathcal{P}_{i,1},\mathcal{P}_{i,2},\cdots,\mathcal{P}_{i,i},-\sigma,-2\sigma,\cdots,-(n-i)\sigma]_{n},$
where $\mathcal{P}_{i,j}$ denotes the original attention score at $(i,j)$, with $\mathcal{P}_{i,i}$ capturing the self-attention along the diagonal. The decay factor $\sigma\cdot(j-i)$ applied for future tokens $i<j$, which enforces causal masking. The standard causal mask operation in Eq. 2 is then modified as:
$\tilde{\mathbf{W}}=\text{SoftMax}(\underbrace{\omega\cdot C+\mathcal{P}}_{\mathbf{W}})\cdot C.$ (9)